# MetaLLM: A High-performant and Cost-efficient Dynamic Framework for Wrapping LLMs

### v2

@@ 4.1 Experimental Setup
We conduct our experiments with LLMs provided by popular API services, including OpenAI and Amazon’s Bedrock. We chose four models from OpenAI: text-ada-001, text-babbage-001, text-curie-001, and text-davinci-002. These models have different costs and different capabilities, giving the users diverse options. The costs of these models are shown in Table 1.
@@
text-ada-001 \$0.0004 / 1K tokens
text-babbage-001 \$0.0005 / 1K tokens
text-curie-001 \$0.0020 / 1K tokens
text-davinci-002 \$0.0200 / 1K tokens
Given a sample SENT, we query OpenAI models with the following prompt:
For Amazon’s Bedrock APIs, we evaluate MetaLLM with four different LLMs: Titan Lite, Cohere Command-Light, Llama 2 7B, and Claude Instant. Since these models come from different families, this evaluation setting exhibits more heterogeneity than in the setting with the OpenAI models. We summarize the cost of each APIs in Table 2.
Table 2: Price of Bedrock APIs.
Model Cost
Titan Lite \$0.00015 / 1K tokens
Cohere Command-Light \$0.00030 / 1K tokens
Llama 2 7B \$0.00075 / 1K tokens
Claude Instant \$0.00080 / 1K tokens
Given a sample SENT, we query Bedrock APIs with the following prompt:
For the paragraph: SENT’, is the sentiment in this paragraph positive or negative? Answer in one word.
@@
We conduct experiments on two text classification datasets: SST-2 and IMDB. SST-2 is a binary sentiment analysis dataset consisting of movie reviews; the task here is to classify whether a review is positive or negative. This dataset has $67,439$ training samples and $872$ test samples; we select $100$ samples from the test set to form the validation set. IMDB dataset has $25,000$ training samples and $25,000$ test samples. We also extract $2,000$ samples from the test set to build the validation set.
@@
Method Test cost Test accuracy
text-ada-001 0.096 80.57
text-babbage-001 0.119 82.12
text-curie-001 0.478 90.54
text-davinci-002 4.779 90.93
MetaLLM $p=0.03$ 0.111 84.02
MetaLLM $p=0.001$ 0.325 90.98
MetaLLM $p=0$ 1.901 91.87
Training MetaLLM. For each input query, we utilize Sentence-BERT [23] model to extract an embedding vector. MetaLLM is a linear model that maps the embedding vector to the reward expectation, which is optimized with the true reward 4 and mean square loss. For a budget $b$, we train MetaLLM with the scaling $p$ five times such that the cost of MetaLLM on the validation set is not higher than $b$ and compute the accuracy of the text classification task. We normalize the cost of each LLM in the reward function such that the highest value is $1$.
@@
Figure 2: The number of samples can be answered by one model but not other models. Cheaper models can answer many queries that more expensive models cannot.
@@ 4.2 Performance of MetaLLM on OpenAI Models
Table 3 shows the accuracy and cost of each OpenAI LLM and MetaLLM with different $p$ on SST-2. As can be observed, the LLM with higher cost has better accuracy. Noticeably, the differences in cost between LLMs can be very high; for example, text-davinci-002 is ten times more expensive than text-curie-001 but only has $0.4\%$ better performance in terms of accuracy. Consequently, querying only the most expensive LLM is not an optimal choice, unless the usage budget of the application is sufficiently high.
Performance of MetaLLM.
For each LLM, except text-ada-001, we compute the cost of choosing that model only, fine-tune the scaling $p$ on the validation set such that the cost is not higher than that budget, train MetaLLM with the found $p$ five times and report the mean accuracy of the text classification task with LLMs selected by MetaLLM. We can observe that MetaLLM  can achieve better accuracy at a lower cost. Specifically, with $p=0.03$, MetaLLM’s performance is ${}2\%$ better than that of defaulting all queries to text-babbage-001, while having a slightly less expensive cost. When $p=0.001$, MetaLLM’s performance is better at that of defaulting to text-curie-001, with a ${}32\%$ cost reduction. More importantly, MetaLLM can reach a higher performance than that of defaulting to text-davinci-002, the best LLM, while spending $60\%$ less.
@@ 4.3 The Heterogeneous Capabilities of Different LLMs
The fact that MetaLLM can achieve better accuracy than even the most expensive OpenAI model means that an LLM’s performance can vary across queries, and cannot always determined by its usage cost (or the size of the LLM). We provide a more rigorous analysis of this observation in this section. In Figure 2, the value at position $(i,j)$ is the number of samples that are correctly classified by the $i$-th model but not the $j$-th model. As we can observe, there are many queries that can be answered by smaller models such as text-ada-001 and text-babbage-001, while text-curie-001 and text-davinci-002 give incorrect answers. If MetaLLM can learn this property and determine the suitable LLM for a query, it will boost the accuracy of using a single model. Our analysis in Section 4.6 confirms this hypothesis.
@@ 4.4 Reward Function With Dynamic Cost
Figure 3: The cost-accuracy trade-off of MetaLLM with the cost in log scale for better visualization. MetaLLM with dynamic cost in the reward function slightly decreases the accuracy with a high budget; however both approaches can perform better than a single LLM.
The cost of an LLM is not fixed for every input; for the zero-shot classification task, it primarily depends on the length of the query. We perform an experiment where we use the exact cost of each training input in the reward function 4 and train MetaLLM with different $p$. The results for this dynamic cost setting and the fixed cost are provided in Figure 3. As we can observe, using dynamic costs does not lead to significant improvements, although both strategies can achieve better accuracies and lower costs than defaulting to a single LLM. When we set a high budget, dynamic cost setting can even has lower accuracy. We hypothesize that using dynamic cost imposes greater penalties on long queries, thereby encouraging MetaLLM to route them to cheaper models; these queries are, however, more complicated and may not be effectively solved by these cheaper models. Therefore, we recommend training MetaLLM with a fixed cost in the reward function for every training query.
4.5 Performance of MetaLLM on Bedrock APIs
This section studies the scenario where the LLMs are heterogeneous, i.e., the less expensive models can perform better on some queries than the more expensive ones. We perform this experiment using the LLMs provided by Amazon’s Bedrock and the IMDB dataset. Table 4 provides the accuracy and the cost of each approach, including defaulting to the same LLM (the first four rows) and our MetaLLM. As we can observe, when defaulting to a single LLM, a more expensive option does not guarantee a better performance; for example, defaulting to Claude Instant, the most expensive LLM, results in lower accuracies compared to other single-LLM options.
On the other hand, training MetaLLM with $p=0$ yields the best performance on the text classification task. Setting a positive value to $p$ decreases the cost substantially while having a minor degradation in the accuracy. Notably, MetaLLM with $p=0.5$ achieves better performance than the best LLM, Llama 2 7B, but only incurs a minimal cost (${}42\%$ less expensive). Compared to the cheapest defaulting option, i.e., Titan Lite, MetaLLMachieves more than 3% better accuracy while a having comparable cost.
Table 4: The cost and accuracy of each LLM on Bedrock APIs
Method Test cost Test accuracy
Titan Lite 0.472 88.10
Cohere Command-Light 0.945 90.65
Llama 2 7B 2.362 91.50
Claude Instant 2.520 87.55
MetaLLM $p=0.5$ 0.485 91.59
MetaLLM $p=0$ 1.376 92.76
@@ 4.6 Analysis of MetaLLM’s Cost Scaling Parameter
Similar observations can be made with Amazon’s Bedrock APIs (Figures 7 and 7). These experiments suggest that even when the application prioritizes performance, putting a small penalty on the cost component can still bring a financial benefit without minimal impact on performance, showing the robustness of using MetaLLM.
@@
Figure 6: The histogram of Bedrock APIs selected by MetaLLM with $p=0$.
Figure 7: The histogram of Bedrock APIs selected by MetaLLM with $p=0.5$.
@@ 5 Conclusion
In this paper, we study the problem of selecting a model in a set of LLMs for optimal performance with cost efficiency. We propose MetaLLM that employs the multi-armed bandit algorithm to learn the reward of querying the correct model with a low price. Our approach is lightweight and applicable to any set of off-the-shelf LLMs and thus versatile in practical use cases. Empirical results show that MetaLLM can improve the accuracy of the best API by around $1\%$ while reducing the cost of zero-shot text classification tasks by $60\%$ and $40\%$ on OpenAI and Bedrock APIs, respectively.

### v3

@@ 4.1 Experimental Setup
We conduct our experiments with LLMs provided by popular API services, including OpenAI and Together AI. We chose four models from OpenAI: text-ada-001, text-babbage-001, text-curie-001, and text-davinci-002. These models have different costs and different capabilities, giving the users diverse options. The costs of these models are shown in Table 2.
Given a sample SENT, we query OpenAI models with the following prompt:
For Together AI APIs, we evaluate MetaLLM with four different LLMs: Gemma, Llama, Mistral, and Qwen. Since these models come from different families, this evaluation setting exhibits more heterogeneity than the setting with the OpenAI models. We summarize the cost of each APIs in Table 2.
Model Cost
text-ada-001 \$0.40 / 1M tokens
text-babbage-001 \$0.50 / 1M tokens
text-curie-001 \$2.00 / 1M tokens
text-davinci-002 \$20.00 / 1M tokens
@@
Gemma \$0.10 / 1M tokens
Llama \$0.18 / 1M tokens
Mistral \$0.20 / 1M tokens
Qwen \$0.30 / 1M tokens
Table 2: Price of Together AI APIs.
Given a sample SENT, we query Together AI APIs with the following prompt:
Answer the following multiple-choice question with only the answer letter (A, B, C, or D):
Question: QUESTION
Options:
A. CHOICE A
B. CHOICE B
C. CHOICE C
D. CHOICE D
Answer:
@@ D. CHOICE D
We conduct experiments on SST-2 – a text classification dataset and MMLU – a question answering dataset. SST-2 is a binary sentiment analysis dataset consisting of movie reviews; the task here is to classify whether a review is positive or negative. This dataset has $67,439$ training samples and $872$ test samples. MMLU is a dataset multitask language understanding dataset including 57 multi-choice question answering subtasks. It has $14,042$ test samples and $99,842$ auxiliary training samples collected from several question answering benchmarks.
Setting Method Test cost Test accuracy
text-ada-001 0.096 80.50
text-babbage-001 0.120 82.80
text-curie-001 0.480 90.60
text-davinci-002 4.800 91.51
Offline MetaLLM (text-babbage-001 budget) 0.120 84.06
MetaLLM (text-curie-001 budget) 0.539 89.56
MetaLLM (text-davinci-002 budget) 2.030 91.97
Online w/o training data MetaLLM (text-babbage-001 budget) 0.303 84.06
MetaLLM (text-curie-001 budget) 0.818 87.27
MetaLLM (text-davinci-002 budget) 1.456 88.30
Online MetaLLM (text-babbage-001 budget) 0.117 84.06
MetaLLM (text-curie-001 budget) 0.455 90.94
MetaLLM (text-davinci-002 budget) 2.046 92.55
@@
Training MetaLLM. For each input query, we utilize Sentence-BERT (Reimers & Gurevych, 2019) model to extract an embedding vector. MetaLLM is a linear model that maps the embedding vector to the reward expectation, which is optimized with the true reward (Eqn. (4)) by Algorithm 1. We normalize the cost of each LLM in the reward function such that the highest value is $1$. For a budget $b$, we train MetaLLM with the scaling $p$ five times, such that the cost of MetaLLM on the validation set is not higher than $b$, and compute the accuracy of the classification task. Given that $p$, we evaluate three different scenarios: i) only optimizing the reward function with training data, ii) only performing online updates during inference, and iii) initializing the policy from training data and updating during inference.
@@ 4.2 Performance of MetaLLM on OpenAI Models
Table 3 shows the accuracy and cost of each OpenAI LLM and MetaLLM with different $p$ on SST-2. As can be observed, the LLM with a higher cost has better accuracy. Noticeably, the differences in cost between LLMs can be very high; for example, text-davinci-002 is ten times more expensive than text-curie-001 but only has $0.4\%$ better performance in terms of accuracy. Consequently, querying only the most expensive LLM is not an optimal choice unless the usage budget of the application is sufficiently high.
Performance of MetaLLM. For each LLM, except text-ada-001, we compute the cost of choosing that model only, fine-tune the scaling $p$ on the validation set such that the cost is not higher than that model’s cost (the budget), train MetaLLM with the found $p$ five times and report the mean accuracy of the text classification task with LLMs selected by MetaLLM. We can observe that MetaLLM  can achieve better accuracy at a lower cost. Specifically, with the same budget as text-babbage-001, MetaLLM’s performance is ${}2\%$ better than that of defaulting all queries to text-babbage-001, while having a slightly less expensive cost. Similarly, MetaLLM can achieve higher accuracy and lower cost than text-curie-001. More importantly, MetaLLM can reach a higher performance than that of defaulting to text-davinci-002, the best LLM, while spending $60\%$ less.
Comparison of different scenarios. Table 3 also demonstrates that only optimizing on training data can yield higher accuracy than text-davinci-002, yet not as effective as online updating on test data. The initialization step also plays an important role; only online updating fails to find a good policy.
Figure 2: The number of samples that can be answered by one model but not by the other models. Cheaper models can answer many queries that more expensive models cannot.
Figure 3: The cost-accuracy trade-off of MetaLLM with the cost in log scale for better visualization. MetaLLM with dynamic cost in the reward function slightly decreases the accuracy with a high budget; however, both approaches can perform better than a single LLM.
@@ 4.3 The Heterogeneous Capabilities of Different LLMs
The fact that MetaLLM can achieve better accuracy than even the most expensive OpenAI model means that an LLM’s performance can vary across queries, and cannot always be determined by its usage cost (or the size of the LLM). In this section, we provide a more rigorous analysis of this observation. In Figure 3, the value at position $(i,j)$ is the number of samples that are correctly classified by the $i$-th model but not the $j$-th model. As we can observe, there exist many queries that can be answered by the smaller models, such as text-ada-001 and text-babbage-001, while text-curie-001 and text-davinci-002 give incorrect answers. If MetaLLM can exploit this heterogeneous relationship between performance and cost to find the most suitable LLM for a query, it can even significantly boost the accuracy of using a single model. Our analysis in Section 4.6 confirms this hypothesis.
@@ 4.4 Reward Function With Dynamic Cost
The cost of an LLM is not fixed for every input; for the zero-shot classification task, it primarily depends on the length of the query. We perform an experiment where we use the exact cost of each training input in the reward function (4) and train MetaLLM with different $p$. The results for this dynamic cost setting and the fixed cost are provided in Figure 3. As we can observe, using dynamic costs does not lead to significant improvements, although both strategies can achieve better accuracies and lower costs than defaulting to a single LLM. When we set a high budget, dynamic cost setting can even has lower accuracy. We hypothesize that using dynamic cost imposes greater penalties on long queries, thereby encouraging MetaLLM to route them to cheaper models; these queries are, however, more complicated and may not be effectively solved by these cheaper models. Therefore, we recommend training MetaLLM with a fixed cost in the reward function for every training query.
4.5 Performance of MetaLLM on Together AI APIs
Setting Method Test cost Test accuracy
Gemma 866.187 64.33
Llama 1559.136 76.69
Mistral 1732.374 66.68
Qwen 2598.561 80.24
Offline MetaLLM (Llama budget) 1474.975 75.70
MetaLLM (Qwen budget) 2287.195 79.68
Online w/o training data MetaLLM (Llama budget) 866.187 64.33
MetaLLM (Qwen budget) 866.187 64.33
Online MetaLLM (Llama budget) 1540.663 76.87
MetaLLM (Qwen budget) 2304.500 80.90
Table 4: The cost and accuracy of each LLM on Together AI APIs
This section studies the scenario where the LLMs are heterogeneous, i.e., the less expensive models can perform better on some queries than the more expensive ones. We perform this experiment using the LLMs provided by Together AI and the MMLU dataset. Table 4 provides the accuracy and the cost of each approach, including defaulting to the same LLM (the first four rows) and our MetaLLM. As we can observe, when defaulting to a single LLM, a more expensive option does not guarantee a better performance; for example, Mistral is more expensive than Llama but yields lower performance.
On the other hand, training MetaLLM with $p=0$ yields the best performance ($80.90\%$ accuracy) on MMLU. Since the distribution of the auxiliary training set is far from the test set, optimizing the reward function does not provide a good policy, i.e., obtaining the highest accuracy. On the other hand, without the initialization step, the multi-arm bandit algorithm only defaults to Gemma. Setting a positive value to $p$ decreases the cost substantially while having a minor degradation in the accuracy. Notably, MetaLLM  can achieve better performance than the second-best LLM, Llama, but with a lower cost. These results further confirm that MetaLLM provides a cost-efficient yet high-performance policy.
@@ 5 Conclusion
In this paper, we study the problem of dynamically selecting an LLM out of a set of LLMs for an input that achieves optimal performance and cost efficiency. To solve this problem, we propose a multi-armed bandit framework that can learn the reward of querying the correct model at a low price. Our approach, denoted as MetaLLM, is lightweight and applicable to any set of off-the-shelf LLMs and thus is versatile in practical use cases. Empirical results show that MetaLLM can improve the accuracy of the best API by around $1\%$ while significantly reducing the cost of zero-shot text classification and multiple-choice question answering tasks by up to $60\%$.